If a 20dB sound level (ticking clock) is replaced by 80dB (garbage disposal), by what factor has the sound intensity been multiplied?

1
Expert's answer
2022-03-06T15:17:50-0500

The sound level is given by

"\beta=10\lg\frac{I}{I_0}"

Hence

"\frac{I_2}{I_1}=10^{\Delta \beta/10}"

"\frac{I_2}{I_1}=10^{60/10}=10^6"